# LLM-Agent-UMF: LLM-based Agent Unified Modeling Framework for Seamless Integration of Multi Active/Passive Core-Agents

### v2

@@ Abstract
In an era where vast amounts of data are collected and processed from diverse sources, there is a growing demand to develop sophisticated AI systems capable of intelligently fusing and analyzing this information. To address these challenges, researchers have turned towards integrating tools into LLM-powered agents to enhance the overall information fusion process. However, the conjunction of these technologies and the proposed enhancements in several state-of-the-art works followed a non-unified software architecture resulting in a lack of modularity and terminological inconsistencies among researchers. To address these issues, we propose a novel LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF) that aims to establish a clear foundation for agent development from both functional and software architectural perspectives. Our framework clearly distinguishes between the different components of an LLM-based agent, setting LLMs, and tools apart from a new element, the core-agent, playing the role of the central coordinator of the agent. This pivotal entity comprises five modules: planning, memory, profile, action, and security—the latter often neglected in previous works. By classifying core-agents into passive and active types based on their authoritative natures, we propose various multi-core agent architectures that combine unique characteristics of distinctive agents to tackle complex tasks more efficiently. We evaluate our framework by applying it to thirteen state-of-the-art agents, thereby demonstrating its alignment with their functionalities and clarifying the overlooked architectural aspects. Moreover, we thoroughly assess five of our proposed architectures through the integration of existing agents into new hybrid active/passive core-agents architectures. This analysis provides clear insights into potential improvements and highlights challenges involved in combining specific agents.
@@
Keywords LLM-based agent  $\cdot$ software architecture  $\cdot$ modularity  $\cdot$ privacy  $\cdot$ security  $\cdot$ classification  $\cdot$ multi-core agent
@@ 1 Introduction
Large Language Models (LLMs) excel in tasks like language modeling, question answering, sentiment analysis, Natural Language Understanding (NLU), commonsense reasoning and knowledge fusion [1, 2]. However, standalone LLMs lacks other skills such as information retrieval, mathematical reasoning, code evaluation, and numerous others. These functional shortcomings can be managed by AI agents leveraging external tools, knowledge repositories and human feedback. An autonomous agent is a system interacting with an environment, sensing it, and acting on it over time following a certain agenda [3]. While there are different classes of agents, in this paper we will be focusing on LLM-based agents which, by combining the capabilities of LLMs and autonomous agents, can achieve a broad range of tasks [1].
@@
Understanding the potential of these agents and improving them necessitates a deep understanding of their structure. In LLM-based agents, besides the LLM who handles reasoning, there are other components that oversee the execution of tasks, ensure the security of the agent, and handle its memory [4]. But merely pinpointing existing functionalities is insufficient for developers and researchers. To provide a comprehensive overview of these agents, survey [5] proposed a comprehensive framework for their construction, consisting of four main modules which will be discussed thoroughly in the next section. Although it provides valuable insights into the functionality of each module in an agent, it overlooks their delineation from a software architectural perspective. This perspective is crucial for developers to establish a common base architecture to build upon and for researchers to improve. Our analysis of the state-of-the-art LLM-based agents reveals common limitations: the complexity of implementation, resulting in unstructured and ambiguous software architecture; lack of modularity and composability, making components non-reusable by other agent-based solutions; and difficulty in maintainability and introducing improvements to existing agents.
In response to these limitations, we propose the LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF). To the best of our knowledge, our framework is the first to emphasize a clear delineation of each component within an LLM-powered agent and define their interactions within specified boundaries from both architectural and functional perspectives. In addition to traditional components like LLMs and tools, we introduce a new unit within the agent, which we label as the "core-agent". This pioneering entity is further classified into two types – active and passive core-agents – which enhance our understanding of each component’s capabilities and accurately describe the dynamics between modules within the agent. As a result, we alleviate the complexity of the architecture and improve the reusability of the components, promoting a shift from multi-agent systems to multi-core agents.
@@
The rest of this paper is organized as follows. Section 2 covers a background on LLM-based agents and reviews relevant state-of-the-art works. Section 3 introduces our LLM-based agent unified modeling framework and possible architectures. Section 4 evaluates the efficiency of our proposed agent designs leveraging the capabilities of multiple distinctive agents. Finally, Section 5 summarizes key findings and discusses future challenges.
@@ 2 Related Work
Figure 1: Framework proposed by survey [5] for LLM-based agents
Existing agents do not necessarily respect these principles as they do not focus primarily on architectural design. This underscores the importance of architectural frameworks as blueprints for LLM-based agents, highlighting their essential role in releasing the full potential of these systems. Such frameworks enable the development of modular, robust, extensible, and interoperable designs. They provide the necessary scaffolding to build increasingly capable and reliable agents capable of tackling complex real-world problems.
@@
3 LLM-based Agent Unified Modeling Framework
In order to comprehensively explore LLM-based agents’ capabilities, a rigorous study was conducted across numerous scholarly databases and online digital libraries. These sources include ScienceDirect, Scopus, Springer, Nature, IEEE Xplore, ACM Digital Library, ACL Anthology, NeurIPS proceedings, AAAI library, arXiv, Google Scholar, Semantic Scholar, among others. The analysis was conducted on articles focusing on tool-powered LLMs, LLM-based agents, LLM planning strategies, information fusion, AI ethical concerns, AI privacy considerations, LLM security and guardrails, software modeling best practices, and terminological consistency in the context of AI systems.
In this section, we start by introducing the fundamental component of our framework, the core-agent, distinguishing it from other elements within an LLM-based agent. By examining the existing literature and analyzing their shortcomings, we delineate a unified approach to designing the internal structure of the core-agent. This comprehensive analysis led us to classify core-agents into two categories: active and passive core-agents. Finally, we outline various multi-core agent architectures that highlight the efficiency of LLM-Agent-UMF in modeling multi-core agent systems.
3.1 Core-Agent: Keystone component of an LLM-based agent
Several works have aimed to establish a well-defined framework for building LLM-based agents. For instance, the paper [16] presents a framework for designing educational problem-solving simulations using LLM-powered agents. The authors emphasize that separating the AI agent from the environment is important in the design process. However, the paper does not provide a clear framework for the agent’s components, making it challenging for future work to identify specific points of modification or reuse. Without a transparent and detailed outline of the agent’s architecture and component interactions, it is difficult to pinpoint where changes can be made to alter the agent’s behavior or how its components could be reused and integrated in other systems.

### v3

# LLM-Agent-UMF: LLM-based Agent Unified Modeling Framework for Seamless Design of Multi Active/Passive Core-Agent Architectures

@@ Abstract
In an era where vast amounts of data are collected and processed from diverse sources, there is a growing demand for sophisticated AI systems capable of intelligently fusing and analyzing this information. To address these challenges, researchers have turned towards integrating tools into LLM-powered agents to enhance the overall information fusion process. However, the conjunction of these technologies and the proposed enhancements in several state-of-the-art works followed a non-unified software architecture, resulting in a lack of modularity and terminological inconsistencies among researchers. To address these issues, we propose a novel LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF) that establishes a clear foundation for agent development from both functional and software architectural perspectives, developed and evaluated using the Architecture Tradeoff and Risk Analysis Framework (ATRAF). Our framework clearly distinguishes between the different components of an LLM-based agent, setting LLMs and tools apart from a new element, the core-agent, which plays the role of central coordinator. This pivotal entity comprises five modules: planning, memory, profile, action, and security—the latter often neglected in previous works. By classifying core-agents into passive and active types based on their authoritative natures, we propose various multi-core agent architectures that combine unique characteristics of distinctive agents to tackle complex tasks more efficiently. We evaluate our framework by applying it to thirteen state-of-the-art agents, thereby demonstrating its alignment with their functionalities and clarifying overlooked architectural aspects. Moreover, we thoroughly assess five architecture variants of our framework by designing new agent architectures that combine characteristics of state-of-the-art agents to address specific goals. Throughout this evaluation, we leveraged ATRAF’s Architectural Framework Tradeoff and Risk Analysis Method (AFTRAM), identifying quality attribute goals, developing scenarios, and analyzing architectural risks, which provided clear insights into potential improvements and highlighted challenges involved in both designing new agents and combining existing ones.
@@
This preprint is the manuscript submitted to and accepted in Information Fusion. The published article is available at https://doi.org/10.1016/j.inffus.2025.103865. This manuscript is available under the CC BY-NC-ND license.
Keywords LLM-based agent $\cdot$ Software architecture $\cdot$ Modularity $\cdot$ Security $\cdot$ Privacy $\cdot$ Safety $\cdot$ Core-agent classification $\cdot$ Multi-core agent
@@ 1 Introduction
Large Language Models (LLMs) excel in tasks like language modeling, question answering, sentiment analysis, Natural Language Understanding (NLU), commonsense reasoning and knowledge fusion [1, 2]. However, standalone LLMs lack other skills such as information retrieval, mathematical reasoning, code evaluation, and numerous others. These functional shortcomings can be managed by AI agents leveraging external tools, knowledge repositories, and human feedback. An autonomous agent is a system interacting with an environment, sensing it, and acting on it over time following a certain agenda [3]. While there are different classes of agents, in this paper we will be focusing on LLM-based agents which, by combining the capabilities of LLMs and autonomous agents, can achieve a broad range of tasks [1].
@@
Understanding the potential of these agents and improving them necessitates a deep understanding of their structure. In LLM-based agents, besides the LLM that handles reasoning, there are other components that oversee the execution of tasks, ensure the security of the agent, and handle its memory [4]. However, merely pinpointing existing functionalities is insufficient for developers and researchers. To provide a comprehensive overview of these agents, the survey [5] proposed a comprehensive framework for their construction, consisting of four main modules which will be discussed thoroughly in the next section. Although it provides valuable insights into the functionality of each module in an agent, it overlooks their delineation from a software architectural perspective. This perspective is crucial for developers to establish a common base architecture to build upon and for researchers to improve. Our analysis of the state-of-the-art LLM-based agents reveals common limitations: the complexity of implementation, resulting in unstructured and ambiguous software architecture; lack of modularity and composability, making components non-reusable by other agent-based solutions; and difficulty in maintainability and introducing improvements to existing agents.
In response to these limitations, we propose the LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF). To the best of our knowledge, our framework is the first to emphasize a clear delineation of each component within an LLM-powered agent and define their interactions within specified boundaries from both architectural and functional perspectives. In addition to traditional components like LLMs and tools, we introduce a new unit within the agent, which we label as the “core-agent”. This pioneering entity is further classified into two types—active and passive core-agents—which enhance our understanding of each component’s capabilities and accurately describe the dynamics between modules within the agent. As a result, we alleviate the complexity of the architecture and improve the reusability of the components, promoting a shift from multi-agent systems to multi-core agents.
@@
The rest of this paper is organized as follows. Section 2 provides a background on LLM-based agents and reviews relevant state-of-the-art works. Section 3 outlines the methodology employed in developing the LLM-Agent-UMF, including the literature review process, quality requirements, design principles, and validation approach. Section 4 introduces our LLM-based Agent Unified Modeling Framework, detailing its components and proposed multi-core agent architectures. Section 5 evaluates the efficiency of our framework by applying it to state-of-the-art agents and designing novel multi-core agent architectures. Finally, Section 6 summarizes key findings and discusses future challenges and directions for enhancing the framework.
@@ 2 Related Work
Figure 1: Framework proposed by survey [5] for LLM-based agents
@@
Existing agents do not consistently adhere to these principles as they do not focus primarily on architectural design. This underscores the importance of architectural frameworks as blueprints for LLM-based agents, highlighting their essential role in releasing the full potential of these systems. Such frameworks enable the development of modular, robust, extensible, and interoperable designs. They provide the necessary scaffolding to build increasingly capable and reliable agents capable of tackling complex real-world problems.
@@
3 Methodology
The Software Development Life Cycle (SDLC) provides a systematic framework for developing robust and scalable systems, encompassing six key phases: Requirements Specification, Software Design, Implementation, Testing, Deployment, and Maintenance. These phases ensure that complex systems, such as LLM-based agents, are designed and delivered with clarity, efficiency, and adaptability. LLM-Agent-UMF focuses primarily on the Requirements Specification and Software Design phases, as these are critical for establishing precise functional and non-functional requirements, crafting architectural blueprints that balance performance, flexibility, and maintainability.
Thoughtful Requirements Specification informs the Software Design phase by defining quality attributes and design principles, which streamline subsequent phases. Modular designs enable efficient Implementation by allowing independent component development, reduce Testing efforts by addressing issues early, facilitate scalable Deployment through strategic architectural choices, and support maintainable architectures for long-term Maintenance. However, to ensure the robustness of the architectural design, a structured evaluation method is essential.
The Architecture Tradeoff and Risk Analysis Framework (ATRAF) [16] provides a robust evaluation method through its Architectural Framework Tradeoff and Risk Analysis Method (AFTRAM), tailored for Architectural Frameworks like the LLM-Agent-UMF, which qualifies as such due to its process-driven methodology, lifecycle support, and cross-domain applicability. AFTRAM’s iterative four-phase spiral process was employed to develop and evaluate LLM-Agent-UMF, with each phase mapped to IMRaD sections per the ATRAF-driven IMRaD Methodology [17] for transparency, rigor, and reproducibility: Phase I (Multi-Context Scenario and Requirements Gathering) elicits scenarios and quality requirements, spanning the Introduction and Methodology sections; Phase II (Architectural Framework Views and Scenario Realization) presents architectural views, detailed in the Methodology section; Phase III (Attribute-Specific Analyses) evaluates quality attributes, covered in the Evaluation and Discussion section; and Phase IV (Sensitivity, Tradeoff, and Risk Analysis) identifies tradeoffs and risks, also in the Evaluation and Discussion section, with interpretations in its Discussion component, presenting a final snapshot of the AFTRAM process.
3.1 Literature Review Process
To develop the LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF), a comprehensive literature review was conducted to identify gaps in existing LLM-based agent frameworks and inform a modular, secure, and scalable design. The review systematically explored scholarly databases and digital libraries, including ScienceDirect, Scopus, Springer, Nature, IEEE Xplore, ACM Digital Library, ACL Anthology, NeurIPS proceedings, AAAI library, arXiv, Google Scholar, and Semantic Scholar. The process targeted literature on tool-augmented LLMs, LLM-based agents, planning strategies, information fusion, AI ethics, privacy, security, software modeling best practices, and terminological consistency in AI systems, ensuring a robust foundation for addressing architectural and functional deficiencies. The review was guided by thirteen specific objectives, each addressing a critical gap or requirement that shaped the framework’s design.
  1. 1.
    Modularity Inspiration from Biological Models: To ensure a modular structure, the review explored cognitive science literature, focusing on human brain modularity, where distinct regions coordinate for cognitive tasks [18]. This informed the application of the Single Responsibility Principle (SRP) to delineate the core-agent’s modules (planning, memory, profile, action, security), ensuring each handles specific functions independently.
  2. 2.
    Structured Analysis for Planning Module: To address the lack of clear component delineation in prior frameworks, the review analyzed LLM-based agent planning literature [19]. This led to a novel four-aspect analysis (process, strategies, techniques, feedback sources) for the planning module, providing a structured approach to task decomposition and plan generation.
  3. 3.
    Human-inspired Task Decomposition: Drawing from cognitive psychology, the review examined human cognitive strategies for breaking down complex tasks [18]. This informed the design of iterative and non-iterative decomposition approaches in the planning module, enhancing its ability to handle complex problems systematically.
  4. 4.
    Criteria-based Technique Selection: To develop a methodology for selecting planning techniques, the review synthesized literature on planning algorithms [19, 20]. Criteria such as task complexity and contextual comprehension were derived, enabling the planning module to choose between rule-based and language model-powered techniques effectively.
  5. 5.
    System Boundary Delineation: To ensure interoperability with external tools and humans, the review studied system architecture literature [21]. This established clear system boundaries for the planning module’s feedback sources (human, tool, sibling core-agent), promoting seamless integration within the agent system.
  6. 6.
    Addressing Terminology Gaps in Memory Module: To resolve unconventional memory terminology in prior works, the review aligned memory scope with human memory models from cognitive science [18]. Terms like short-term and long-term memory were adopted, replacing ambiguous Unified and Hybrid Memory terms, enhancing clarity and consistency.
  7. 7.
    Architectural Relevance for Memory Location: Prioritizing software architecture relevance, the review analyzed software engineering literature to focus on memory location (embedded vs. extension) [22]. This approach addressed gaps in prior frameworks that emphasized scope over architectural considerations, improving memory management design.
  8. 8.
    Categorization of Memory Formats: To clarify memory representation, the review synthesized literature on data representation formats (natural language, embeddings, SQL databases, structured lists) [23, 24]. This led to the novel three-perspective analysis (scope, location, format) for the memory module, addressing literature gaps in representation clarity.
  9. 9.
    Best Practice Alignment for Security Module: To ensure the security module adhered to modern standards, the review examined industry best practices and AI security standards [25, 26]. The Secure by Design principle was adopted, integrating security from the outset to address high-risk AI deployment challenges.
  10. 10.
    Advancement Over C.I.A. Triad Usage: To overcome limitations in prior works using the C.I.A. triad (Confidentiality, Integrity, Availability) as a risk-classification schema, the review analyzed cybersecurity literature [27]. The triad was integrated as a foundational architectural principle, guiding the security module’s design for robust protection.
  11. 11.
    Comprehensive Security Research: To address multifaceted security challenges, the review explored AI security literature covering asset protection, threat defense, incident mitigation, and privacy compliance [13, 28]. This informed the security module’s layered safeguarding mechanisms (prompt, response, data privacy safeguarding).
  12. 12.
    Guardrail Methodology Analysis: To operationalize security measures, the review analyzed guardrail methodologies from LLM service providers and open-source communities [28, 29]. Rule-based and LLM-based guardrails were identified, forming the security module’s operational layer for effective threat mitigation.
  13. 13.
    Regulatory Compliance for Trust Boundaries: To ensure compliance with data protection obligations, the review examined data protection regulations [25]. Transparent trust boundaries were architected for the security module, ensuring privacy and user confidence in interactions with external tools.
The literature review process was iterative, with findings cross-referenced to validate gaps and ensure alignment with software engineering and AI development best practices. By addressing modularity, terminology, security, and architectural clarity, the review provided a solid foundation for the LLM-Agent-UMF, enabling it to overcome the limitations of prior frameworks and establish a unified, extensible, and secure approach to LLM-based agent design.
3.2 Quality Requirements
The LLM-Agent-UMF was designed to meet a comprehensive set of non-functional quality requirements to ensure robust, efficient, and trustworthy LLM-based agent systems. These requirements, derived from an analysis of existing frameworks and software engineering best practices, address modularity, security, scalability, and other critical attributes that enhance the framework’s applicability and the performance of agents designed using it. The following quality requirements were explicitly addressed in the framework’s design, while others such as recovery were considered for future research.
  • Flexibility: Allows the framework to adapt to varying requirements, environments, or use cases, ensuring agents can handle diverse tasks and dynamic contexts effectively.
  • Extensibility: Enables the framework to be expanded with new features or capabilities, supporting the integration of additional modules or functionalities without disrupting existing components.
  • Reusability: Facilitates the use of framework components across different agents or applications, reducing duplication and enhancing development efficiency.
  • Modifiability: Enables straightforward alterations or updates to accommodate changes, maintaining system relevance and functionality over time.
  • Maintainability: Ensures that the framework and its agents can be easily updated or fixed, supporting long-term system evolution and operational continuity.
  • Scalability: Supports handling increased loads or growth, allowing agents to manage expanding tasks or user demands without performance degradation.
  • Dynamic Adaptability: Allows agents to adjust to changing conditions or requirements at runtime, improving responsiveness and versatility.
  • Resilience and Fault Tolerance: Ensures agents can withstand disruptions and continue operating despite failures or errors, enhancing reliability and robustness in unpredictable environments.
  • Performance: Optimizes operational efficiency, including speed, responsiveness, and resource usage, to deliver high-quality user experiences.
  • Availability: Ensures reliable system operation and accessibility, minimizing downtime and maintaining service reliability for users.
  • Interoperability: Facilitates seamless interaction with other systems or components, enabling effective information exchange and integration.
  • Security: Protects against unauthorized access, threats, or data breaches, incorporating integrity to maintain the accuracy, completeness, and uncorrupted state of data, processes, and components, thereby safeguarding system reliability and user confidence.
  • Privacy: Protects personal or sensitive data from unauthorized access or disclosure, ensuring compliance with data protection standards and fostering user trust.
  • Safety and Ethical Alignment: Ensures adherence to ethical standards and principles, promoting user trust and societal acceptance of agent actions.
3.3 Design Principles
The LLM-Agent-UMF is grounded in a set of software engineering design principles that guide its development and serve as recommendations for architects and designers creating LLM-based agents based on the framework. These principles, derived from software engineering best practices and tailored to address the challenges of LLM-based agent systems, ensure modularity, security, scalability, and clarity. They support the quality requirements outlined in Section 3.2, such as Flexibility, Maintainability, and Security, by providing actionable strategies for architectural design. The principles are organized into three categories—general architectural principles, security and privacy principles, and framework-level design principles—each contributing to the framework’s robustness and applicability.
General architectural principles:
  • Terminological Consistency: Uniform use of terms and concepts ensures Maintainability and Modifiability by reducing ambiguity. The lack of standardized terminology across researchers and practitioners has led to discrepancies in describing LLM-based agent components, making this a primary concern for clear communication and collaboration [30].
  • Modularity, Single Responsibility Principle (SRP), and Separation of Concerns: Systems are broken into smaller, independent modules, each with a single responsibility and distinct functionality, to enhance Maintainability, Scalability, Reusability, Modifiability, and Performance. This combined principle ensures clear delineation of responsibilities, reducing functional overlap [21].
  • Open-Closed Principle (OCP) and Loose Coupling: Systems are open for extension but closed for modification, with components designed to have minimal dependencies, supporting Modifiability, Maintainability, Scalability, Flexibility, and Extensibility. This principle enables new functionalities without altering existing components and ensures independent updates [31].
  • Composability: Components are designed to be combined or reused to build larger systems, supporting Flexibility, Scalability, Maintainability, Reusability, and Extensibility, enabling versatile system configurations.
Security and privacy principles:
  • Security by Design (SbD): Security is integrated into the development process from the outset to achieve Security and Integrity, ensuring robust protection against threats [27, 32].
  • Privacy by Design (PbD): Privacy considerations are integrated from the start to achieve Privacy, protecting sensitive data and aligning with data protection standards [33, 34].
  • C.I.A. Triad: The Confidentiality, Integrity, and Availability (C.I.A.) triad serves as a foundational principle to ensure secure and reliable systems, safeguarding data confidentiality, maintaining data and process integrity, and ensuring system availability [27].
  • Trust Boundaries Definition: Zones of trust are clearly defined to manage security and access control, supporting Security, Integrity, and Privacy, ensuring compliance with data protection requirements [28].
  • Defense in Depth: Multiple layers of security controls are implemented to protect against threats, achieving Security and Integrity, enhancing system resilience [35].
Framework-level design principles:
  • Technology-Agnostic Design: The framework avoids dependency on specific technologies to promote Flexibility, Maintainability, and Interoperability, allowing architects and designers to select technologies best suited for their LLM-based agents while ensuring the framework’s flexibility to support these choices.
  • Variability Management: Variations in requirements or configurations are handled systematically, supporting Flexibility, Extensibility, Maintainability, and Reusability, accommodating diverse agent designs.
  • Extensibility Mechanisms: Extension points, such as pluggable modules and support for multiple LLMs, enable future enhancements without core changes, supporting Extensibility, Maintainability, Flexibility, and Reusability.
  • Constraint-based Design: Rules and boundaries are established to guide the framework’s design, ensuring consistent and predictable behavior, supporting Performance, Maintainability, Safety, Security, Integrity, and Availability.
These design principles collectively ensure that the LLM-Agent-UMF is robust, adaptable, and secure, while providing a clear blueprint for designing LLM-based agents. By adhering to these principles, architects can create agents that meet the quality requirements outlined in Section 3.2, fostering Modularity, Scalability, and Trustworthiness in complex AI systems.
3.4 Contributions: Core-Agent as Keystone Component
Prior frameworks for LLM-based agents, such as those proposed by [5, 7], and [36], have primarily focused on the functional perspective, detailing capabilities like planning, memory, and action. However, these works often neglect issues from a software architectural perspective, failing to delineate the boundaries of different software components within the agent. This lack of architectural clarity results in ambiguous designs, reduced Modularity, and challenges in Reusability and Maintainability. LLM-Agent-UMF addresses these limitations by clearly delineating software components, introducing the core-agent as the keystone component that serves as the central coordinator, as depicted in Figure 2.